### v1

@@ 5.5 Evaluation on Distributed Inference
The debloating results of distributed inference with 8 GPUs are still as significant as those of single-GPU inference. For inference with 8 GPUs, vLLM uses slightly fewer shared libraries than the results of a single GPU, while Transformers uses nearly the same number of libraries. The reductions in file size, host code size, function count, and device code size align closely with the results using a single GPU. However, the element count reduction in device code is lower than that of single-GPU inference, suggesting that distributed inference utilizes more kernels in device code. Additionally, the reduction metrics are similar across different models, once again indicating that different models do not significantly affect the bloat in the ML frameworks.
Summary. Debloating results of vLLM and Transformers using distributed inference on 8 GPUs are consistent with those of single-GPU inference.
5.6 Kernel Detector Performance Overhead
@@
To assess the performance overhead introduced by the kernel detector on the target ML workload, we first run the PyTorch Training MobileNetV2 workload 10 times and recorded the average execution time. Next, we ran the same workload, this time with tracing enabled by the kernel detector and by NSys respectively, each for 10 runs. We then compare the average execution times across the three setups.
The average execution time for the original workload was 180 seconds. With the kernel detector enabled, the execution time increased to 253 seconds, a 41% overhead. In contrast, tracing with NSys increases the execution time to 407 seconds, introducing a 126% overhead. The kernel detector imposes significantly lower overhead than NSys, making it a more practical choice for detecting used kernels, especially for long-running workloads like ML training. We note that this is a one time overhead that can occur in, e.g., a profiling stage before the actual deployment.
@@ 6 Discussion
As Sculley et al. Sculley et al. (2015) highlight, ML systems have hidden technical debt due to a set of ML-specific issues. Our work emphasizes this technical debt by exposing the hidden bloat in ML frameworks. Software bloat has been a long-standing issue in the software industry. While existing debloating research has focused on traditional software, bloat in ML systems is not well understood. The uniqueness of ML systems is that they contain both host and device code, leading to significant bloat. Our evaluations over four ML frameworks across ten workloads with around 300 shared libraries show significant bloat in these frameworks, with up to 72% size reduction in host code and 75% size reduction in device code. Furthermore, unlike traditional software, ML frameworks experience substantial bloat in device code. All this bloat leads to increased storage needs, higher memory usage, and longer execution times.
In our evaluations, only a small subset of code is consistently utilized. This suggests that code unused by one workload is likely unnecessary for others as well. Besides, existing research focuses solely on unused code, overlooking “used bloat” — code executed by a workload but not contributing meaningfully to the performance or functionality. For instance, when training a model with a specific optimizer, the optimizer may initialize a context with numerous non-essential function calls. Compared with PyTorch, the larger host code size but smaller reduction in TensorFlow may indicate the presence of “used bloat”. Such “used bloat” is particularly harmful, as it is executed and thus consumes scarce memory, CPU, and GPU resources. Moreover, it is more difficult to detect as it is actually executed. Future research could focus on identifying and eliminating this “used bloat”.
Shared libraries in ML frameworks are significantly larger than traditional shared libraries Zhang et al. (2024), mainly due to device code as shown in this work. As storage and network bandwidth are critical bottlenecks in edge data centers Richins et al. (2021), the file size reduction achieved by debloating can help alleviate these bottlenecks.
@@ 7 Conclusion
We propose a novel approach to debloat ML frameworks by removing unnecessary code in both host and device code within shared libraries. Our approach first detects kernel names used in device code by running ML workloads, such as training or inference a model. We then locate the used kernels within the shared libraries and remove the unused code. We implement our approach based on an existing debloating tool, and evaluate it on four ML frameworks across ten workloads over 300 shared libraries. Our evaluation shows that ML frameworks are highly bloated, with up to 55% size reduction in shared libraries in these frameworks. We show that up to 75% of device code and 99% of device elements being unnecessary for target ML workloads; Up to 72% of host code and 93% of host functions are unnecessary for target ML workloads. This bloat not only increases storage overhead but also degrades runtime performance. Debloating these frameworks reduces the peak memory usage, peak GPU memory usage, and startup time by up to 74.6%, 69.6%, and 44.6%, respectively.
@@ Appendix A Appendix
Framework Model #Lib. Total File Size/MB Host Code Device Code
@@
nsys profile trace=cuda -o report
{Command to run the ML workload}

### v3

@@ 4.5 Evaluation on Different GPUs
Next, we present debloating results for additional ML workloads using vLLM and Transformers. We select the top 9 LLMs from the Hugging Face Open LLM Leaderboard Beeching et al. (2023) and deploy them on both frameworks using distributed inference with 8×A100 40GB GPUs. This setup aims to evaluate whether debloating works with distributed inference under a different hardware setup. Results are detailed in Table 10 in the Appendix. The debloating results of distributed inference with 8 GPUs are still as significant as those of single-GPU inference. For inference with 8 GPUs, vLLM uses slightly fewer shared libraries than the results of a single GPU, while Transformers uses nearly the same number of libraries. The reductions in file size, CPU code size, function count, and GPU code size align closely with the results using a single GPU. However, the element count reduction in GPU code is lower than that of single-GPU inference, suggesting that distributed inference utilizes more kernels in GPU code. Additionally, the reduction metrics are similar across different models, once again indicating that different models do not significantly affect the bloat in the ML frameworks.
Summary. Negativa-ML effectively debloats ML shared libraries across different GPUs and distributed inference setups.
4.6 Performance of Negativa-ML
In this section, we evaluate the performance of Negativa-ML in terms of the time taken to debloat shared libraries. We first evaluate the end-to-end time, i.e., the time taken from running the target ML workload until obtaining the debloated shared libraries. Then, we specifically evaluate the overhead introduced to the target ML workload by the kernel detector.
Table 8 shows the end-to-end time taken by Negativa-ML to debloat shared libraries for each workload. The time varies depending on three factors: the original execution time of the workload, the number of shared libraries involved, and the number of functions (kernels) utilized by the workload. We note that the times taken are to debloat all the libraries in the workload. If only debloating a single library, the time needed will be significantly shorter. Moreover, the debloating process is a one-time overhead that can occur in, e.g., a preprocessing stage prior to deployment.
Table 8: End-to-end time taken by Negativa-ML to debloat shared libraries for each workload.
Model Framework Operation #Lib. Time/s
MobileNetV2 PyTorch Train 113 651
Inference 111 383
TensorFlow Train 253 659
Inference 251 585
Transformer PyTorch Train 154 1,247
Inference 154 344
TensorFlow Train 398 18,420
Inference 396 639
Llama2 vLLM Inference 170 713
Transformers Inference 98 362
@@ 4.6 Performance of Negativa-ML
To assess the performance overhead introduced by the kernel detector on the target ML workload, we first run the PyTorch Training MobileNetV2 workload 10 times and recorded the average execution time. Next, we ran the same workload, this time with tracing enabled by the kernel detector and by NSys respectively, each for 10 runs. We then compare the average execution times across the three setups. The average execution time for the original workload was 180 seconds. With the kernel detector enabled, the execution time increased to 253 seconds, a 41% overhead. In contrast, tracing with NSys increases the execution time to 407 seconds, introducing a 126% overhead. The kernel detector imposes significantly lower overhead than NSys, making it a more practical choice for detecting used kernels, especially for long-running workloads like ML training.
Summary. Debloating is a one-time overhead that can be performed in a preprocessing stage prior to deployment. The kernel detector introduces a 41% overhead for the first run, significantly lower than traditional tracing tools like NSys which introduce a 126% overhead.
@@ 5 Discussion
As Sculley et al. Sculley et al. (2015) highlight, ML systems have hidden technical debt due to a set of ML-specific issues. Our work emphasizes this technical debt by exposing the hidden bloat in ML frameworks. Software bloat has been a long-standing issue in the software industry. While existing debloating research has focused on traditional software, bloat in ML systems is not well understood. The uniqueness of ML systems is that they contain both CPU and GPU code, leading to significant bloat. Our evaluations over four ML frameworks across ten workloads with around 300 shared libraries show significant bloat in these frameworks, with up to 72% size reduction in CPU code and 75% size reduction in GPU code. Furthermore, unlike traditional software, ML frameworks experience substantial bloat in GPU code. All this bloat leads to increased storage needs, higher memory usage, and longer execution times.
In our evaluations, only a small subset of code is consistently utilized. This suggests that code unused by one workload is likely unnecessary for others as well. Besides, existing research focuses solely on unused code, overlooking “used bloat” — code executed by a workload but not contributing meaningfully to the performance or functionality. For instance, when training a model with a specific optimizer, the optimizer may initialize a context with numerous non-essential function calls. Compared with PyTorch, the larger CPU code size but smaller reduction in TensorFlow may indicate the presence of “used bloat”. Such “used bloat” is particularly harmful, as it is executed and thus consumes scarce memory, CPU, and GPU resources. Moreover, it is more difficult to detect as it is actually executed. Future research could focus on identifying and eliminating this “used bloat”.
Shared libraries in ML frameworks are significantly larger than traditional shared libraries Zhang et al. (2024), mainly due to GPU code as shown in this work. As storage and network bandwidth are critical bottlenecks in edge data centers Richins et al. (2021), the file size reduction achieved by debloating can help alleviate these bottlenecks.
@@ 6 Conclusion
We propose a novel approach to debloat ML frameworks by removing unnecessary code in both CPU and GPU code within shared libraries. Our approach first detects kernel names used in GPU code by running ML workloads, such as training or inference a model. We then locate the used kernels within the shared libraries and remove the unused code. We implement our approach based on an existing debloating tool, and evaluate it on four ML frameworks across ten workloads over 300 shared libraries. Our evaluation shows that ML frameworks are highly bloated, with up to 55% size reduction in shared libraries in these frameworks. We show that up to 75% of GPU code and 99% of GPU elements being unnecessary for target ML workloads; Up to 72% of CPU code and 93% of CPU functions are unnecessary for target ML workloads. This bloat not only increases storage overhead but also degrades runtime performance. Debloating these frameworks reduces the peak memory usage, peak GPU memory usage, and startup time by up to 74.6%, 69.6%, and 44.6%, respectively.
Acknowledgment This work is funded by a WASP PhD student grant and an SSF grant (grant #FFL21-0091).
@@ Appendix A Appendix
Framework Model #Lib. Total File Size/MB CPU Code GPU Code
@@
nsys profile --trace=cuda -o report {Command to run the ML workload}